# Low-Altitude ISAC with Rotatable Active and Passive Arrays

### v1

@@ III-B3 Closed form solution for the QCQP
With $\{\mu_{k},y_{k}\}$ fixed and $g_{a}$ replaced by $\mathcal{U}_{a}$, (P2) reduces to the convex quadratic constrained quadratic program (QCQP) stated as below
$\text{(P2.1):}\underset{\boldsymbol{W}}{\max}\,$ $\sum_{k=1}^{K}{\mathcal{Y}\left(\boldsymbol{W}\mid\mu_{k},y_{k}\right)}-\rho\max\{\mathcal{U}_{a}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)}),0\}$ (41)
s.t. $\left(22\text{c}\right).$
In general, (P2.1) can be solved using convex optimization tools. However, we derive an efficient solution based on Lagrange duality and the Karush-Kuhn-Tucker (KKT) conditions. Specifically, if the MSE threshold is violated at the current $\boldsymbol{W}^{(t)}$, the objective function (41) can be compactly rewritten as a summation with a linear and a quadratic term
$\underset{\boldsymbol{W}}{\max}\,$ $-\text{tr}(\boldsymbol{W}^{H}\boldsymbol{QW})\,\,+2\text{Re}\{\text{Tr}(\boldsymbol{P}^{H}\boldsymbol{W})\}\quad\text{s.t.}\,\,\text{(22c)},$ (42)
where $\boldsymbol{Q}$ and $\boldsymbol{P}$ are defined by
@@
$\boldsymbol{P}=\boldsymbol{P}_{C}+\rho\boldsymbol{P}_{\text{S}},$ (43b)
@@
$\boldsymbol{Q}_{C}=\frac{1}{\ln 2}\sum_{k=1}^{K}{\left(1+\mu_{k}\right)\left|y_{k}\right|^{2}\boldsymbol{f}_{k}\boldsymbol{f}_{k}^{H}},$ (44a)
$\boldsymbol{Q}_{\text{S}}=\frac{1}{2A}\sum_{a=1}^{A}{L_{a}^{\text{lip}}}\boldsymbol{I}_{M+K},$ (44b)
$\boldsymbol{P}_{C}=\frac{1}{\ln 2}\left[\left(1+\mu_{1}\right)y_{1}\boldsymbol{f}_{1},\dots,\left(1+\mu_{K}\right)y_{K}\boldsymbol{f}_{K}\right],$ (44c)
$\boldsymbol{P}_{\text{S}}=\frac{1}{2A}\sum_{a=1}^{A}{\boldsymbol{G}_{a}\left(\boldsymbol{W}^{\left(t\right)}\right)+L_{a}^{\text{lip}}\boldsymbol{W}^{\left(t\right)}}.$ (44d)
@@
$\mathcal{L}(\boldsymbol{W},\nu)=$ $-\text{tr}(\boldsymbol{W}^{H}\boldsymbol{QW})+2\text{Re}\{\text{tr}(\boldsymbol{P}^{H}\boldsymbol{W})t\}$
@@
$-\boldsymbol{QW}^{*}+\boldsymbol{P}-\nu^{*}\boldsymbol{I}=0,$ (46a)
@@
where $\nu$ is found by bisection search to satisfy the power constraint (22c) with equality. Since $\nu$ enters as a scalar shift on the identity, we precompute the eigenvalue decomposition (EVD) $\boldsymbol{Q}=\boldsymbol{U\varLambda U}^{H}$ with $\boldsymbol{\varLambda}=\text{diag}(\lambda_{1},\dots,\lambda_{M})$ and $\lambda_{m}$ being the eigenvalue; thus, for $\nu>-\min\{\lambda_{m}\}_{m=1}^{M}$, we have
@@
where $\cdot]_{i,:}$ denote the $i$-th row of a matrix. Furthermore, if the MSE threshold is satisfied in current $\boldsymbol{W}^{(t)}$, we let $\boldsymbol{Q}=\boldsymbol{Q}_{\text{C}}$ and $\boldsymbol{P}=\boldsymbol{P}_{C}$ and perform the same procedures in (43)-(49).
@@ III-B4 Overall algorithm and complexity analysis for subproblem 1
The overall low-complexity algorithm for solving (P2) is summarized in Algorithm 1. Each inner iteration performs one EVD of an $M{\times}M$ Hermitian matrix at cost $\mathcal{O}(M^{3})$ and a scalar bisection whose per-step cost is $\mathcal{O}(M(K+M))$ to calculate $\|\boldsymbol{W}(\nu)\|_{F}^{2}$; the number of bisection steps grows only logarithmically with precision. The MM construction provides a tight lower bound of the penalized objective; therefore, the penalized objective in (P2) is non-decreasing over $\boldsymbol{W}$-updates. Under the power ball, the sequence $\{\boldsymbol{W}^{(t)}\}$ admits limit points that satisfy the first-order stationarity conditions of the surrogate built from MM and QT.
@@
1:  Input: $\{\boldsymbol{f}_{k}\}_{k=1}^{K}$, $\{\boldsymbol{f}_{\mathrm{S},a}\}_{a=1}^{A}$, $\{\mathcal{P}_{\mathrm{d},a}\}$, $P_{\mathrm{B}}$, $\rho$, $\eta_{\mathrm{th}}$.
@@
4:   Compute $\|\boldsymbol{S}_{a}\|_{2}^{2}$ via (27), $L_{a}^{\text{lip}}$ via (30), $\boldsymbol{G}_{a}(\boldsymbol{W}^{(t)})$ via (28), and $\mathcal{U}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)})$ via (29).
5:   Build $\boldsymbol{Q}_{\mathrm{C}},\boldsymbol{P}_{\mathrm{C}}$ from (44); if $\mathcal{U}(\boldsymbol{W}^{(t)}\mid\boldsymbol{W}^{(t)})>\eta_{\mathrm{th}}$, also build $\boldsymbol{Q}_{\mathrm{S}},\boldsymbol{P}_{\mathrm{S}}$, else set them to zero. Set $\boldsymbol{Q}$ and $\boldsymbol{P}$ via (43).
6:   Compute EVD $\boldsymbol{Q}=\boldsymbol{U}\boldsymbol{\Lambda}\boldsymbol{U}^{H}$. Find $\nu^{\star}$ by bisection using (45); set $\boldsymbol{W}^{(t{+}1)}=\boldsymbol{W}(\nu^{\star})$ via (48).
7:   Update auxiliaries $\mu_{k}^{(t{+}1)}$ by (38) and $y_{k}^{(t{+}1)}$ by(39).
@@ III-C Subproblem 2: Passive Beamforming
$\text{(P3)}:\underset{\boldsymbol{\theta}}{\max}\,\,$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}(\boldsymbol{\theta})\right)}$
$-\rho\max\left\{\text{MSE}(\boldsymbol{\theta})-\eta_{\text{th}},0\right\}$ (50)
s.t. $\left(22\text{d}\right).$
The unique challenge in RIS passive beamforming is the unit-modulus constraint (22d). Nevertheless, it inherently defines a Riemannian manifold, known as the complex circle manifold
@@ III-C1 Problem reformulation
where $\boldsymbol{\Xi}_{\text{S},a}$, $\boldsymbol{\xi}_{\text{S},a}$, and $\xi_{\text{S},a}$ are computed based on (14) by
$\!\boldsymbol{\Xi}_{\mathrm{S},a}=G^{\mathrm{R}}(\boldsymbol{r}^{\mathrm{R}})\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}})\boldsymbol{B}^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}})^{H},\!\!$ (53a)
$\!\boldsymbol{\xi}_{\mathrm{S},a}^{H}=\!\sqrt{G^{\mathrm{B}}(\boldsymbol{r}^{\mathrm{B}})G^{\mathrm{R}}(\boldsymbol{r}^{\mathrm{R}})}(\boldsymbol{t}_{a}^{\mathrm{BT}})^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}}),\!\!\!$ (53b)
$\!\xi_{\mathrm{S},a}=G^{\mathrm{B}}(\boldsymbol{r}^{\mathrm{B}})(\boldsymbol{t}_{a}^{\mathrm{BT}})^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{t}_{a}^{\mathrm{BT}}.\!\!$ (53c)
@@
$\!\!\!S_{a}(\boldsymbol{\theta})$ $=\boldsymbol{\theta}^{H}\boldsymbol{\Xi}_{\mathrm{S},a}\boldsymbol{\theta}+2\Re\{\boldsymbol{\xi}_{\mathrm{S},a}^{H}\boldsymbol{\theta}\}+\xi_{\mathrm{S},a}-\mathcal{P}_{\mathrm{d},a},$ (56a)
@@
$F\left(\boldsymbol{\theta}\right)=$ $\sum_{k=1}^{K}{\log_{2}}\Big(1+\frac{C_{1,k}\left(\boldsymbol{\theta}\right)}{C_{2,k}\left(\boldsymbol{\theta}\right)}\Big)$
$-\rho\max\Big\{\frac{1}{A}\sum_{a=1}^{A}{|S_{a}\left(\boldsymbol{\theta}\right)|^{2}}-\eta_{\text{th}},0\Big\},$ (57)
@@ III-C2 RCG method
To solve (P3.1), we adopt the Riemannian conjugate gradient (RCG) method, which generalizes the classical conjugate gradient algorithm over curved spaces, as a manifold is a topological space that locally resembles the Euclidean space [28],[29]. At any point, tangent vectors give all feasible updating directions, and their set is the tangent space
@@
Using the chain rule, the complete Euclidean gradient expression $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta})$ can be computed by (62) at the top of the next page, which is suitable for the case that the MSE threshold does not satisfy in the current $\boldsymbol{\theta}^{(\jmath)}$; otherwise, omit the MSE-related terms in (62).
$\nabla_{\boldsymbol{\theta}}F\left(\boldsymbol{\theta}\right)=\frac{1}{\ln 2}\sum_{k=1}^{K}{\frac{C_{2,k}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}C_{1,k}\left(\boldsymbol{\theta}\right)-C_{1,k}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}C_{2,k}\left(\boldsymbol{\theta}\right)}{C_{2,k}\left(\boldsymbol{\theta}\right)\left(C_{1,k}\left(\boldsymbol{\theta}\right)+C_{2,k}\left(\boldsymbol{\theta}\right)\right)}}-\frac{\rho}{A}\sum_{a=1}^{A}{2}S_{a}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}S_{a}\left(\boldsymbol{\theta}\right)$ (62a)
$\qquad\,\,\,\quad=\frac{2}{\ln 2}\sum_{k=1}^{K}{\frac{C_{2,k}\left(\mathbf{\Xi}_{\text{C},k,k}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{C},k,k}\right)-C_{1,k}\sum_{i=1,i\neq k}^{K}{\mathbf{\Xi}_{\text{C},k,i}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{C},k,i}}}{C_{2,k}\left(C_{1,k}+C_{2,k}\right)}}-\frac{4\rho}{A}\sum_{a=1}^{A}{S_{a}}\left(\boldsymbol{\theta}\right)\left(\mathbf{\Xi}_{\text{S},a}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{S},a}\right)$ (62b)
@@
$\boldsymbol{\tau}^{(\jmath)}=-\nabla_{\mathcal{R}_{\boldsymbol{\theta}}}F^{\left(\jmath\right)}+\beta_{\boldsymbol{\theta}}^{(\jmath)}\mathsf{Proj}_{\boldsymbol{\theta}^{(\jmath)}}(\boldsymbol{\tau}^{(\jmath-1)}),$ (65)
@@
4:  Compute $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta}^{(\jmath)})$ via (62) and $\mathsf{Rgrad}F(\boldsymbol{\theta})$ via (63).

### v2

@@ III-B3 Closed form solution for the QCQP
With $\{\mu_{k},\eta_{k}\}$ fixed and $g_{a}$ replaced by $\mathcal{U}_{a}$, (P2) reduces to the convex quadratic constrained quadratic program (QCQP) stated as below
$\text{(P2.1):}\underset{\boldsymbol{W}}{\max}\,$ $\sum_{k=1}^{K}{\mathcal{Y}\left(\boldsymbol{W}\!\mid\!\mu_{k},\eta_{k}\right)}-\frac{\rho}{D}\sum_{a=1}^{A}{\mathcal{U}_{a}(\boldsymbol{W}\!\mid\!\boldsymbol{W}^{(t)})}$ (42)
[…]
In general, (P2.1) can be solved using convex optimization tools. However, we derive a solution based on Lagrange duality and the Karush-Kuhn-Tucker (KKT) conditions. Specifically, at the current $\boldsymbol{W}^{(t)}$, the objective (42) can be compactly rewritten as a summation with a linear and a quadratic term
$\underset{\boldsymbol{W}}{\max}\,\,$ $-\mathrm{Tr}(\boldsymbol{W}^{H}\boldsymbol{QW})+2\text{Re}\{\mathrm{Tr}(\boldsymbol{P}^{H}\boldsymbol{W})\}$ (43)
[…]
where we omit constant values that do not affect optimization, and the matrices $\boldsymbol{Q}$ and $\boldsymbol{P}$ are defined respectively by
@@
$\boldsymbol{P}=\boldsymbol{P}_{\text{C}}+\rho\boldsymbol{P}_{\text{S}},$ (44b)
@@
$\boldsymbol{Q}_{\text{C}}=\frac{1}{\ln 2}\sum_{k=1}^{K}{\left(1+\mu_{k}\right)\left|\eta_{k}\right|^{2}\boldsymbol{f}_{k}\boldsymbol{f}_{k}^{H}},$ (45a)
$\boldsymbol{Q}_{\text{S}}=\frac{1}{2D}\sum_{a=1}^{A}{L_{a}^{\text{lip}}}\boldsymbol{I}_{M},$ (45b)
$\boldsymbol{P}_{\text{C}}=\frac{1}{\ln 2}\left[\left(1+\mu_{1}\right)\eta_{1}\boldsymbol{f}_{1},\dots,\left(1+\mu_{K}\right)\eta_{K}\boldsymbol{f}_{K}\right],$ (45c)
$\boldsymbol{P}_{\text{S}}=\frac{1}{2D}\sum_{a=1}^{A}{L_{a}^{\text{lip}}\boldsymbol{W}^{\left(t\right)}-\boldsymbol{G}_{a}\big(\boldsymbol{W}^{(t)}\big)}.$ (45d)
@@
$\mathcal{L}(\boldsymbol{W},\nu)=$ $-\mathrm{Tr}(\boldsymbol{W}^{H}\boldsymbol{QW})+2\text{Re}\{\mathrm{Tr}(\boldsymbol{P}^{H}\boldsymbol{W})\}$
@@
$-(\boldsymbol{Q}+\nu^{*}\boldsymbol{I})\boldsymbol{W}^{*}+\boldsymbol{P}=0,$ (47a)
@@
where $\nu$ is found by bisection search to satisfy the power constraint (24b) with equality. Since $\nu$ enters as a scalar shift on the identity, we precompute the eigenvalue decomposition (EVD) $\boldsymbol{Q}=\boldsymbol{U\varLambda U}^{H}$ with $\boldsymbol{\varLambda}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{M})$ and $\lambda_{m}$ being the eigenvalue; thus, for $\nu>-\min\{\lambda_{m}\}_{m=1}^{M}$, we have
@@
where $\cdot]_{i,:}$ denote the $i$-th row of a matrix.
@@ III-B4 Overall algorithm and complexity analysis for subproblem 1
The overall low-complexity algorithm for solving (P2) is summarized in Algorithm 1. Each inner iteration performs one EVD of an $M{\times}M$ Hermitian matrix at cost $\mathcal{O}(M^{3})$ and a scalar bisection whose per-step cost is $\mathcal{O}(M(K+M))$ to calculate $\|\boldsymbol{W}(\nu)\|_{F}^{2}$; the number of bisection steps grows only logarithmically with precision. The MM construction provides a tight lower bound of the objective in (P2), which is non-decreasing over $\boldsymbol{W}$-updates. Under the power ball, the sequence $\{\boldsymbol{W}^{(t)}\}$ admits limit points that satisfy the first-order stationarity conditions of the surrogate built from MM and QT.
@@
1:  Input: $\{\boldsymbol{f}_{k}\}_{k=1}^{K}$, $\{\boldsymbol{f}_{\text{S},a}\}_{a=1}^{A}$, $\{\mathcal{P}_{\text{d},a}\}$, $P_{\text{B}}$, $\rho$.
@@
4:   Compute $\|\boldsymbol{S}_{a}\|_{2}^{2}$ via (28), $L_{a}^{\text{lip}}$ via (31), $\boldsymbol{G}_{a}(\boldsymbol{W}^{(t)})$ via (29a), and $\mathcal{U}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)})$ via (30).
5:   Build $\boldsymbol{Q}_{\text{C}}$, $\boldsymbol{P}_{\text{C}}$, $\boldsymbol{Q}_{\text{S}}$, and $\boldsymbol{P}_{\text{S}}$ from (45). Set $\boldsymbol{Q}$ and $\boldsymbol{P}$ via (44).
6:   Perform EVD by $\boldsymbol{Q}=\boldsymbol{U}\boldsymbol{\Lambda}\boldsymbol{U}^{H}$. Find $\nu^{\star}$ by bisection using (49) and (50); set $\boldsymbol{W}^{(t{+}1)}=\boldsymbol{W}(\nu^{\star})$ via (48).
7:   Update auxiliaries $\mu_{k}^{(t{+}1)}$ by (39) and $\eta_{k}^{(t{+}1)}$ by (40).
@@ III-C Subproblem 2: Passive Beamforming
$\text{(P3)}:\underset{\boldsymbol{\theta}}{\max}\,\,$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}(\boldsymbol{\theta})\right)}-\rho\text{NMSE}(\boldsymbol{\theta})$ (51)
[…]
The unique challenge in RIS passive beamforming is the unit-modulus constraint. Nevertheless, it inherently defines a Riemannian manifold, known as the complex circle manifold
@@ III-C1 Problem reformulation
where $\boldsymbol{\Xi}_{\text{S},a}$, $\boldsymbol{\xi}_{\text{S},a}$, and $\xi_{\text{S},a}$ are computed based on (II-C) by
$\!\boldsymbol{\Xi}_{\mathrm{S},a}=G^{\text{R}}(\boldsymbol{r}^{\text{R}})\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}})\boldsymbol{B}^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}})^{H},\!\!$ (54a)
$\!\boldsymbol{\xi}_{\mathrm{S},a}^{H}=\!\sqrt{G^{\text{B}}(\boldsymbol{r}^{\text{B}})G^{\text{R}}(\boldsymbol{r}^{\text{R}})}(\boldsymbol{t}_{a}^{\mathrm{BT}})^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}}),\!\!\!$ (54b)
$\!\xi_{\mathrm{S},a}=G^{\text{B}}(\boldsymbol{r}^{\text{B}})(\boldsymbol{t}_{a}^{\mathrm{BT}})^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{t}_{a}^{\mathrm{BT}}.\!\!$ (54c)
@@
$\!\!\!S_{a}(\boldsymbol{\theta})$ $=\boldsymbol{\theta}^{H}\boldsymbol{\Xi}_{\mathrm{S},a}\boldsymbol{\theta}+2\Re\{\boldsymbol{\xi}_{\mathrm{S},a}^{H}\boldsymbol{\theta}\}+\xi_{\mathrm{S},a}-\bar{\mathcal{P}}_{\mathrm{d},a},$ (57a)
@@
$F(\boldsymbol{\theta})=\sum_{k=1}^{K}\log_{2}\!\left(1+\frac{C_{1,k}(\boldsymbol{\theta})}{C_{2,k}(\boldsymbol{\theta})}\right)-\frac{\rho}{D}\sum_{a=1}^{A}\bigl|S_{a}(\boldsymbol{\theta})\bigr|^{2},$ (58)
@@ III-C2 RCG method
To solve (P3.1), we adopt the RCG method, which generalizes the classical conjugate gradient algorithm over curved spaces, as a manifold is a topological space that locally resembles the Euclidean space [38, 39]. At any point, tangent vectors give all feasible updating directions, and their set is the tangent space
@@
Using the chain rule, the complete Euclidean gradient expression $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta})$ can be computed by (63) at the top of the next page.
$\nabla_{\boldsymbol{\theta}}F\left(\boldsymbol{\theta}\right)=\frac{1}{\ln 2}\sum_{k=1}^{K}{\frac{C_{2,k}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}C_{1,k}\left(\boldsymbol{\theta}\right)-C_{1,k}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}C_{2,k}\left(\boldsymbol{\theta}\right)}{C_{2,k}\left(\boldsymbol{\theta}\right)\left(C_{1,k}\left(\boldsymbol{\theta}\right)+C_{2,k}\left(\boldsymbol{\theta}\right)\right)}}-\frac{\rho}{D}\sum_{a=1}^{A}{2}S_{a}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}S_{a}\left(\boldsymbol{\theta}\right)$ (63a)
$\qquad\,\,\,\quad=\frac{2}{\ln 2}\sum_{k=1}^{K}{\frac{C_{2,k}\left(\mathbf{\Xi}_{\text{C},k,k}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{C},k,k}\right)-C_{1,k}\sum_{i=1,i\neq k}^{K}{\mathbf{\Xi}_{\text{C},k,i}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{C},k,i}}}{C_{2,k}\left(C_{1,k}+C_{2,k}\right)}}-\frac{4\rho}{D}\sum_{a=1}^{A}{S_{a}}\left(\boldsymbol{\theta}\right)\left(\mathbf{\Xi}_{\text{S},a}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{S},a}\right)$ (63b)
@@
$\boldsymbol{\tau}^{(\jmath)}=-\nabla_{\mathcal{R}_{\boldsymbol{\theta}}}F^{\left(\jmath\right)}+\chi^{(\jmath)}\mathsf{Proj}_{\boldsymbol{\theta}^{(\jmath)}}(\boldsymbol{\tau}^{(\jmath-1)}),$ (66)
@@
4:  Compute $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta}^{(\jmath)})$ via (63)) and $\mathsf{Rgrad}F(\boldsymbol{\theta})$ via (64).